# Are Your Models Still Fair? Fairness Attacks on Graph Neural Networks via Node Injections

### v1

@@ 4.2 Node injection with principles
Uncertainty-maximization principle. Intuitively, nodes with higher model uncertainties are positioned closer to the decision boundary, which means their predicted labels are more vulnerable and easier to flip when facing adversarial attacks. We acknowledge that the model uncertainty may not be the only method to measure the vulnerability of nodes, and we will discuss potential alternative approaches in Appendix H.4. Inspired by (Liu et al., 2022), we utilize a Bayesian GNN to estimate the model uncertainty of each node, where we employ the Monte Carlo dropout approach (Gal and Ghahramani, 2016) to approximate the distributions of the sampled model parameters. Given a GNN with parameters $\theta_{\mathcal{B}}$, we obtain different model parameters through $T$ times independent Bernoulli dropout sampling processes, i.e.:
$\begin{array}[]{r@{}c@{}l}P(M_{i})&\sim&\mbox{Bernoulli}(p)\\[2.0pt] \theta_{\mathcal{B}_{i}}&\enspace=&M_{i}\odot\theta_{\mathcal{B}}\end{array},i\in\{1,2,\dots,T\}$ (8)
where $M_{i}$ is the $i$th sampled binary mask following the Bernoulli distribution with parameter $p$, and $\odot$ denotes the dot production operation. Here we take a two-layer GCN (Hamilton et al., 2017) with parameters $\theta_{\mathcal{B}}$ as the Bayesian GNN for estimating the uncertainty of each node, and $\theta_{\mathcal{B}}$ is optimized by minimizing the following objective function, which consists of a cross-entropy loss plus a regularization term:
$L(\theta_{\mathcal{B}})=-\frac{1}{T}\sum_{i=1}^{T}{\mathcal{Y}\log(f_{\theta_{\mathcal{B}_{i}}}(\mathcal{V},\mathcal{G}))}+\frac{1-p}{2T}\|\theta_{\mathcal{B}}\|_{2}^{2}$ (9)
@@
$U=\mathop{Var}_{i=1}^{T}(f_{\theta_{\mathcal{B}_{i}}}(\mathcal{V},\mathcal{G}))$ (10)
@@
Intuitively, compared with random node injection, our strategy prevents information propagation between nodes of different sensitive groups through the injected nodes, making it easier to accentuate differences in embeddings between groups and thereby exacerbate unfairness issues (Wang et al., 2022a). We also provide a brief theoretical analysis to show that such a strategy could lead to the increase of node-level homophily ratio. Similar to (Du et al., 2022), we define the node-level homophily-ratio $\mathcal{H}_{u}$ as the ratio of neighbors of node $u$ that have the same sensitive attribute as node $u$, i.e. $\mathcal{H}_{u}=\frac{\sum_{v\in\mathcal{N}_{u}}\mathbbm{1}(s_{u}=s_{v})}{|\mathcal{N}_{u}|}$, where $\mathcal{N}_{u}$ denotes the neighbors of node $u$ and $\mathbbm{1}(\cdot)$ is an indicator function. Then we can have:
@@
Due to space limitation, the proof for Lemma 1 is provided in Appendix C. It is worth noting that $\mathcal{H}_{u}$ is also equivalent to the probability of choosing neighbors with the same sensitive attribute for node $u$. From the perspective of message propagation, higher node-level homophily-ratio indicates that more sensitive-related information will be aggregated to the target node, thus leading to more severe unfairness issues on sensitive attributes [^2] Similar conclusions have been concluded from multiple prior works (Luo et al., 2023; Mehrabi et al., 2021; Yang et al., 2024).. We believe that such characteristics could empower our node injection strategy with stronger capability on fairness attacks.
@@ 4.3 Feature optimization
Classification loss. Although our primary goal is to maximize the unfairness of a GNN model, it is crucial to ensure that the utility of the victim model will not experience a significant decrease after training on a poisoned graph (Hussain et al., 2022; Kang et al., 2023; Zhang et al., 2024), thus being unnoticeable for utility-based attack detection. To this end, we set cross-entropy loss as our first objective function, i.e.:
$L_{CE}=-\frac{1}{|\mathcal{V}^{tr}|}\sum_{u\in\mathcal{V}^{tr}}{y_{u}\log h_{u}}$ (11)
@@
$L_{SP}=-\|\frac{1}{|\mathcal{V}^{tr}_{0}|}\sum_{u\in\mathcal{V}^{tr}_{0}}{h_{u}}-\frac{1}{|\mathcal{V}^{tr}_{1}|}\sum_{u\in\mathcal{V}^{tr}_{1}}{h_{u}}\|^{2}_{2}$ (12)
$L_{EO}=-\|\sum_{y\in\mathcal{Y}}(\frac{1}{|\mathcal{V}_{0,y}^{tr}|}\sum_{u\in\mathcal{V}_{0,y}^{tr}}{h_{u,y}}-\frac{1}{|\mathcal{V}_{1,y}^{tr}|}\sum_{u\in\mathcal{V}_{1,y}^{tr}}{h_{u,y}})\|^{2}_{2}$ (13)
@@
$L_{CF}=-\|\frac{1}{|\mathcal{V}_{I,0}|}\sum_{u\in\mathcal{V}_{I,0}}{\textbf{X}_{I,u}}-\frac{1}{|\mathcal{V}_{I,1}|}\sum_{u\in\mathcal{V}_{I,1}}{\textbf{X}_{I,u}}\|^{2}_{2}$ (14)
@@
$L=L_{CE}+\alpha\cdot L_{CF}+\beta\cdot(L_{SP}+L_{EO})$ (15)
@@ 5.1 Experimental settings
Datasets. Experiments are conducted on three real-world datasets namely Pokec-z, Pokec-n, and DBLP. Both Pokec-z and Pokec-n are subgraphs sampled from Pokec, one of the largest online social networks in Slovakia, according to the provinces of users (Dai and Wang, 2021). Each node in these graphs represents a user, while each edge represents an unidirectional following relationship. The datasets provide node attributes including age, gender, and hobbies, and the classification task is to predict the working fields of users. DBLP is a coauthor network dataset (Hussain et al., 2022), where each node represents an author and two authors will be connected if they publish at least one paper together. The node features are constructed based on the words selected from the corresponding author’s published papers. The final classification task is to predict the research area of the authors. The detailed dataset statistics are summarized in Table 1.
Victim models. As a gray-box attack method, we target multiple classical GNNs as victim models, including GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), APPNP (Klicpera et al., 2019), and SGC (Wu et al., 2019). We also include three well-established fairness-aware GNNs — FairGNN (Dai and Wang, 2021), FairVGNN (Wang et al., 2022b) and FairSIN (Yang et al., 2024) as our selected victim models. The details of these victim models will be elaborated in Appendix E.
@@
Implementation details. As shown in Table 1, only a part of the nodes have the label information, and we randomly select 50%, 25%, and 25% labeled nodes as the training set, validation set, and test set, respectively. In line with the prior work (Hussain et al., 2022), we choose region as the sensitive attribute for Pokec-z and Pokec-n, and gender for DBLP. For all victim models, we employ a two-layer GCN model as the surrogate model. Due to space limitations, please refer to Appendix G for more reproducibility details. Our code and data are released at: https://github.com/LuoZhhh/NIFA.
@@ 5.2 Main attack performance
To comprehensively evaluate NIFA’s effectiveness, we employ multiple mainstream GNNs including GCN, GraphSAGE, APPNP and SGC, besides three classical fairness-aware GNNs namely FairGNN, FairVGNN and FairSIN as our victim models. We record the average accuracy, $\Delta_{SP}$ and $\Delta_{EO}$ before and after conducting our poisoning attack on the victim models five times. The experimental results are reported in Table 2, and we can have the following observations:
@@
  • On three fairness-aware models, FairGNN, FairVGNN and FairSIN, NIFA still causes noticeable fairness impacts. For example, the $\Delta_{EO}$ of FairVGNN on Pokec-z increases from 2.59% to 9.28%, a nearly fourfold increase. It indicates that even fairness-aware GNN models are also vulnerable to our attack, highlighting the urgency of proposing more robust fairness mechanisms.
@@ 6 Defense discussion to fairness attacks on GNNs
Reliable training nodes. One key assumption in NIFA is that the nodes with high model uncertainty will be much easier to be attacked, which has also been verified by the ablation study in Appendix H.2. In this way, administrators can pay more attention to these nodes and their abnormal neighbors for defense purposes. For example, engineers can pre-train a model to detect the abnormal nodes or edges in advance, especially those that emerged recently in the training data, and weaken their impacts on the model by randomly masking these nodes or edges in the input graph or decreasing their weights in the message propagation during the training of GNNs.
Figure 2: Defense performance on Pokec-z with masking $\eta$ training nodes with the highest uncertainty.
@@
To verify our assumption, we conduct a simple experiment by removing a proportion of nodes ($\eta$) with the highest uncertainty $U$ from supervision signals after the attack. Similarly, GCN (Hamilton et al., 2017) is employed as the victim model and we gradually tune $\eta$ from 0 to 0.6 with step 0.1, where $\eta=0$ means no defense is involved. The performance of NIFA on the Pokec-z dataset with different $\eta$ is illustrated in Figure 2. It can be seen that, since NIFA mainly focuses on attacking nodes with high uncertainty, after masking a part of these nodes during the training stage, the fairness attack performance of NIFA gradually decreases with a small fluctuation in accuracy. However, it is worth noting that although such an intuitive strategy can defend the attack from NIFA to some extent, there is still obvious fairness deterioration compared with the performance of clean GCN in Table 2 ($\Delta_{SP}$=7.13, $\Delta_{EO}$=5.10). More dedicated and effective defense mechanisms in the future are still in demand.
@@
Strengthen the connections among groups. One main reason behind the success of NIFA in fairness attack is the guidance of the homophily-increase principle during node injection. The ablation study in Appendix H.2 also provides empirical evidence for this claim. As we analyze in Section 4.2, NIFA will lead to the increase of node-level homophily-ratio, which means more sensitive-related information will be aggregated and enlarged within the group. Given this, we believe that an effective defensive strategy is to strengthen the information propagation among different sensitive groups, thus preventing the risks of information cocoons (Luo et al., 2023; Masrour et al., 2020) and fairness issues.
@@
Fairness auditing. At last, we find that a crucial assumption in NIFA and other research (Hussain et al., 2022; Kang et al., 2023; Zhang et al., 2024) is that GNN model administrators will only audit the utility metrics of the models, such as accuracy or F1-scores. Therefore, as long as attackers can ensure that the model utility is not affected excessively, it will be hard for administrators to realize the attack. Consequently, we strongly suggest that model administrators should also incorporate fairness-related metrics into their monitoring scopes, especially before model deployment or during the beta testing phase, thus, mitigating the potential broader negative impacts and social risks. For instance, if an updated GNN model suddenly demonstrates obvious fairness deterioration compared with the previous versions, the model administrators should be careful about the potential fairness attacks. However, the challenge of this approach mainly lies in the diverse definitions of fairness, such as group fairness (Dai and Wang, 2021; Luo et al., 2022), individual fairness (Dong et al., 2021), etc., and group fairness based on different sensitive attributes (Dai and Wang, 2021; Wang et al., 2022b) or structures (Liu et al., 2023; Luo et al., 2023) may further lead to different definitions. Therefore, model administrators might need prior knowledge or expertise to determine what kinds of fairness metrics to be included in their monitoring scopes.
@@ 7 Conclusion
Limitations. Firstly, NIFA is still under the settings of gray-box attacks, which requires accessibility to the labels and sensitive attributes. We acknowledge that such information may not always be available and we leave the extensions to the more realistic black-box attack settings as future work. Moreover, although we present some insights on the defense strategies of GNN fairness, more effective defense measures are still under-explored, calling for more future research efforts.
@@ Appendix B Attack settings
Attack stage. Attacks can be categorized into two types according to the time when the attacks take place [Zhang et al., 2022]: poisoning attack and evasion attack. Poisoning attacks occur at the training phase of victim models, which will lead to poisoned models. In contrast, evasion attacks target the inference phase, and can not affect the model parameters. In this work, NIFA belongs to the poisoning attacks.
Attacker’s knowledge. Generally, according to the knowledge of attackers, the attack methods can be categorized into three types including white-box attack, black-box attack and gray-box attack [Zhang et al., 2022]. As we introduced in Section 3, we propose NIFA within the gray-box attack settings to make our attack more practical in the real world, which is also consistent with multiple prior research on GNN attacks [Hussain et al., 2022, Sun et al., 2020, Zhang et al., 2024]. Different from white-box attacks and black-box attacks, gray-box attacks mean that attackers can only access the training data, including the input graph $\mathcal{G}$, the labels $\mathcal{Y}$ and the sensitive attribute $s$ of each node. Note that, the model architecture and parameters are invisible to attackers under the gray-box attack settings, which leads to that the attackers need to train a surrogate model in advance to assess the effectiveness of their proposed attacks.
Attacker’s capability. One merit of NIFA is that there is no need for the attackers to have the authority to modify the existing graph structure, such as adding or deleting edges between existing real nodes, or modifying the existing real nodes’ features. In contrast, NIFA injects $\mathcal{V}_{I}$ malicious nodes and poisons the graph $\mathcal{G}$ into $\mathcal{G}^{\prime}$ to launch an attack, which is much more practical for the attackers in the real world. For example, in social networks the attackers only need to sign up multiple zombie accounts and interact with real accounts. Note that, different from some injection-based attack [Ju et al., 2023, Sun et al., 2020], NIFA will not modify the training set and true node label set $\mathcal{Y}$, as such operations are typically infeasible in the real world. The intrinsic idea of NIFA is to impact the GNN training process through massage propagation on a poisoned graph.
Within the gray-box attack settings, we also assume that the attackers have sufficient computational resources and budget to train a surrogate model and have access to the real graph as input. Similar to prior attack methods [Sun et al., 2020], attackers are also required to set thresholds $b$ and $d$ for the number of injected nodes and their degrees respectively to make NIFA deceptive and unnoticeable.

### v2

@@ 4.2 Node injection with principles
Uncertainty-maximization principle. Intuitively, nodes with higher model uncertainties are positioned closer to the decision boundary, which means their predicted labels are more vulnerable and easier to flip when facing adversarial attacks. We acknowledge that the model uncertainty may not be the only method to measure the vulnerability of nodes, and we will discuss potential alternative approaches in Appendix H.2. Inspired by [20], we utilize a Bayesian GNN to estimate the model uncertainty of each node, where we employ the Monte Carlo dropout approach [9] to approximate the distributions of the sampled model parameters. Given a GNN with parameters $\theta_{\mathcal{B}}$, we obtain different model parameters through $T$ times independent Bernoulli dropout sampling processes, i.e.:
$\begin{array}[]{r@{}c@{}l}P(M_{i})&\sim&\mbox{Bernoulli}(p)\\[2.0pt] \theta_{\mathcal{B}_{i}}&\enspace=&M_{i}\odot\theta_{\mathcal{B}}\end{array},i\in\{1,2,\dots,T\},$ (8)
where $M_{i}$ is the $i$th sampled binary mask following the Bernoulli distribution with parameter $p$, and $\odot$ denotes the dot production operation. Here we take a two-layer GCN [10] with parameters $\theta_{\mathcal{B}}$ as the Bayesian GNN for estimating the uncertainty of each node, and $\theta_{\mathcal{B}}$ is optimized by minimizing the following objective function, which consists of a cross-entropy loss plus a regularization term:
$L(\theta_{\mathcal{B}})=-\frac{1}{T}\sum_{i=1}^{T}{\mathcal{Y}\log(f_{\theta_{\mathcal{B}_{i}}}(\mathcal{V},\mathcal{G}))}+\frac{1-p}{2T}\|\theta_{\mathcal{B}}\|_{2}^{2},$ (9)
@@
$U=\mathop{Var}_{i=1}^{T}(f_{\theta_{\mathcal{B}_{i}}}(\mathcal{V},\mathcal{G})).$ (10)
@@
Intuitively, compared with random node injection, our strategy prevents information propagation between nodes of different sensitive groups through the injected nodes, making it easier to accentuate differences in embeddings between groups and thereby exacerbate unfairness issues [37]. We also provide a brief theoretical analysis to show that such a strategy could lead to the increase of node-level homophily ratio. Similar to [8], we define the node-level homophily-ratio $\mathcal{H}_{u}$ as the ratio of neighbors of node $u$ that have the same sensitive attribute as node $u$, i.e. $\mathcal{H}_{u}=\frac{\sum_{v\in\mathcal{N}_{u}}\mathbbm{1}(s_{u}=s_{v})}{|\mathcal{N}_{u}|}$, where $\mathcal{N}_{u}$ denotes the neighbors of node $u$ and $\mathbbm{1}(\cdot)$ is an indicator function. Then we can have:
@@
Due to space limitation, the proof for Lemma 1 is provided in Appendix C. It is worth noting that $\mathcal{H}_{u}$ is also equivalent to the probability of choosing neighbors with the same sensitive attribute for node $u$. From the perspective of message propagation, higher node-level homophily-ratio indicates that more sensitive-related information will be aggregated to the target node, thus leading to more severe unfairness issues on sensitive attributes [^2] Similar conclusions have been concluded from multiple prior works [22, 26, 40]. For better a understanding of the relationship between the homophily-ratio and unfairness, one can also refer to [40], which provides the corresponding theoretical analysis from a massage propagation perspective.. We believe that such characteristics could empower our node injection strategy with stronger capability on fairness attacks.
@@ 4.3 Feature optimization
Classification loss. Although our primary goal is to maximize the unfairness of a GNN model, it is crucial to ensure that the utility of the victim model will not experience a significant decrease after training on a poisoned graph [11, 13, 41], thus being unnoticeable for utility-based attack detection. To this end, we set cross-entropy loss as our first objective function, i.e.:
$L_{CE}=-\frac{1}{|\mathcal{V}^{tr}|}\sum_{u\in\mathcal{V}^{tr}}{y_{u}\log h_{u}},$ (11)
@@
$L_{SP}=-\|\frac{1}{|\mathcal{V}^{tr}_{0}|}\sum_{u\in\mathcal{V}^{tr}_{0}}{h_{u}}-\frac{1}{|\mathcal{V}^{tr}_{1}|}\sum_{u\in\mathcal{V}^{tr}_{1}}{h_{u}}\|^{2}_{2},$ (12)
$L_{EO}=-\|\sum_{y\in\mathcal{Y}}(\frac{1}{|\mathcal{V}_{0,y}^{tr}|}\sum_{u\in\mathcal{V}_{0,y}^{tr}}{h_{u,y}}-\frac{1}{|\mathcal{V}_{1,y}^{tr}|}\sum_{u\in\mathcal{V}_{1,y}^{tr}}{h_{u,y}})\|^{2}_{2},$ (13)
@@
$L_{CF}=-\|\frac{1}{|\mathcal{V}_{I,0}|}\sum_{u\in\mathcal{V}_{I,0}}{\textbf{X}_{I,u}}-\frac{1}{|\mathcal{V}_{I,1}|}\sum_{u\in\mathcal{V}_{I,1}}{\textbf{X}_{I,u}}\|^{2}_{2},$ (14)
@@
$L=L_{CE}+\alpha\cdot L_{CF}+\beta\cdot(L_{SP}+L_{EO}),$ (15)
@@ 5.1 Experimental settings
Datasets. Experiments are conducted on three real-world datasets namely Pokec-z, Pokec-n, and DBLP. Both Pokec-z and Pokec-n are subgraphs sampled from Pokec, one of the largest online social networks in Slovakia, according to the provinces of users [5]. Each node in these graphs represents a user, while each edge represents an unidirectional following relationship. The datasets provide node attributes including age, gender, and hobbies, and the classification task is to predict the working fields of users. DBLP is a coauthor network dataset [11], where each node represents an author and two authors will be connected if they publish at least one paper together. The node features are constructed based on the words selected from the corresponding author’s published papers. The final classification task is to predict the research area of the authors. The detailed dataset statistics are summarized in Table 1.
Victim models. As a gray-box attack method, we target multiple classical GNNs as victim models, including GCN [14], GraphSAGE [10], APPNP [15], and SGC [39]. We also include three well-established fairness-aware GNNs — FairGNN [5], FairVGNN [38], and FairSIN [40] as our selected victim models. The details of these victim models will be elaborated in Appendix E.
@@
Implementation details. As shown in Table 1, only a part of the nodes have the label information, and we randomly select 50%, 25%, and 25% labeled nodes as the training set, validation set, and test set, respectively. In line with the prior work, [11], we choose region as the sensitive attribute for Pokec-z and Pokec-n, and gender for DBLP. For all victim models, we employ a two-layer GCN model as the surrogate model. Due to space limitations, please refer to Appendix G for more reproducibility details.
@@ 5.2 Main attack performance
To comprehensively evaluate NIFA’s effectiveness, we employ multiple mainstream GNNs including GCN, GraphSAGE, APPNP, and SGC, besides three classical fairness-aware GNNs namely FairGNN, FairVGNN, and FairSIN as our victim models. We record the average accuracy, $\Delta_{SP}$ and $\Delta_{EO}$ before and after conducting our poisoning attack on the victim models five times. The experimental results are reported in Table 2, and we can have the following observations:
@@
  • On three fairness-aware models, FairGNN, FairVGNN, and FairSIN, NIFA still causes noticeable fairness impacts. For example, the $\Delta_{EO}$ of FairVGNN on Pokec-z increases from 2.59% to 9.28%, a nearly fourfold increase. It indicates that even fairness-aware GNN models are also vulnerable to our attack, highlighting the urgency of proposing more robust fairness mechanisms.
@@ 5.3 Comparison with other attack Models
5.4 Ablation study
Figure 2: Ablation study of each module in NIFA
The results are reported in Figure 2, where we can have the following observations. Firstly, after removing the uncertainty-maximization principle (NIFA-U), the fairness attack performance consistently decreases on three datasets. This is expected since the concept of uncertainty helps NIFA find more vulnerable nodes, thus improving the attack effectiveness. Secondly, after removing the homophily-increase principle (NIFA-H), the attack performance drops obviously, which verifies the homophily-ratio is crucial in GNN fairness. Finally, without iterative training during feature optimization (NIFA-I), the attack performance decreases slightly on all datasets. The main reason is that the iterative training strategy could help NIFA to have better robustness to dynamic victim models.
@@ 6 Defense discussion to fairness attacks on GNNs
Figure 3: Defense performance on Pokec-z with masking $\eta$ training nodes with the highest uncertainty
@@
Reliable training nodes. One key assumption in NIFA is that the nodes with high model uncertainty will be much easier to be attacked, which can also be supported by the ablation study in Section 5.4. In this way, administrators can pay more attention to these nodes and their abnormal neighbors for defense purposes. For example, engineers can pre-train a model to detect the abnormal nodes or edges in advance, especially those that emerged recently in the training data, and weaken their impacts on the model by randomly masking these nodes or edges in the input graph or decreasing their weights in the message propagation during the training of GNNs.
@@
To verify our assumption, we conduct a simple experiment by removing a proportion of nodes ($\eta$) with the highest uncertainty $U$ from supervision signals after the attack. Similarly, GCN [10] is employed as the victim model and we gradually tune $\eta$ from 0 to 0.6 with step 0.1, where $\eta=0$ means no defense is involved. The performance of NIFA on the Pokec-z dataset with different $\eta$ is illustrated in Figure 3. It can be seen that, since NIFA mainly focuses on attacking nodes with high uncertainty, after masking a part of these nodes during the training stage, the fairness attack performance of NIFA gradually decreases with a small fluctuation in accuracy. However, it is worth noting that although such an intuitive strategy can defend the attack from NIFA to some extent, there is still obvious fairness deterioration compared with the performance of clean GCN in Table 2 ($\Delta_{SP}$=7.13, $\Delta_{EO}$=5.10). More dedicated and effective defense mechanisms in the future are still in demand.
@@
Strengthen the connections among groups. One main reason behind the success of NIFA in fairness attack is the guidance of the homophily-increase principle during node injection. The ablation study in Appendix 5.4 also provides empirical evidence for this claim. As we analyze in Section 4.2, NIFA will lead to the increase of node-level homophily-ratio, which means more sensitive-related information will be aggregated and enlarged within the group. Given this, we believe that an effective defensive strategy is to strengthen the information propagation among different sensitive groups, thus preventing the risks of information cocoons [22, 25] and fairness issues.
@@
Fairness auditing. At last, we find that a crucial assumption in NIFA and other research [11, 13, 41] is that GNN model administrators will only audit the utility metrics of the models, such as accuracy or F1-scores. Therefore, as long as attackers can ensure that the model utility is not affected excessively, it will be hard for administrators to realize the attack. Consequently, we strongly suggest that model administrators should also incorporate fairness-related metrics into their monitoring scopes, especially before model deployment or during the beta testing phase, thus, mitigating the potential broader negative impacts and social risks. For instance, if an updated GNN model suddenly demonstrates obvious fairness deterioration compared with the previous versions, the model administrators should be careful about the potential fairness attacks. However, the challenge of this approach mainly lies in the diverse definitions of fairness, such as group fairness [5, 23], individual fairness [6], etc., and group fairness based on different sensitive attributes [5, 38] or structures [21, 22] may further lead to different definitions. Therefore, model administrators might need prior knowledge or expertise to determine what kinds of fairness metrics to be included in their monitoring scopes.
@@ 7 Conclusion
Limitations. Firstly, NIFA is still under the settings of gray-box attacks, which requires accessibility to the labels and sensitive attributes. We acknowledge that such information may not always be available and we leave the extensions to the more realistic black-box attack settings as future work. Moreover, although we present some insights on the defense strategies of GNN fairness, more effective defense measures are still under-explored, calling for more future research efforts. At last, currently we only focus on fairness based on sensitive attributes, while neglecting the fairness based on graph structures. Since different fairness may stem from different sources, we leave this as our future work.
Acknowledgement
The work is supported by the National Natural Science Foundation of China (No.62172174). The authors would also like to thank Xiran Song and Jianxun Lian for their suggestions and contributions.
@@ Appendix B Attack settings
Attack stage. Attacks can be categorized into two types according to the time when the attacks take place [42]: poisoning attack and evasion attack. Poisoning attacks occur at the training phase of victim models, which will lead to poisoned models. In contrast, evasion attacks target the inference phase, and can not affect the model parameters. In this work, NIFA belongs to the poisoning attacks.
Attacker’s knowledge. Generally, according to the knowledge of attackers, the attack methods can be categorized into three types including white-box attack, black-box attack and gray-box attack [42]. As we introduced in Section 3, we propose NIFA within the gray-box attack settings to make our attack more practical in the real world, which is also consistent with multiple prior research on GNN attacks [11, 32, 41]. Different from white-box attacks and black-box attacks, gray-box attacks mean that attackers can only access the training data, including the input graph $\mathcal{G}$, the labels $\mathcal{Y}$ and the sensitive attribute $s$ of each node. Note that, the model architecture and parameters are invisible to attackers under the gray-box attack settings, which leads to that the attackers need to train a surrogate model in advance to assess the effectiveness of their proposed attacks.
Attacker’s capability. One merit of NIFA is that there is no need for the attackers to have the authority to modify the existing graph structure, such as adding or deleting edges between existing real nodes, or modifying the existing real nodes’ features. In contrast, NIFA injects $\mathcal{V}_{I}$ malicious nodes and poisons the graph $\mathcal{G}$ into $\mathcal{G}^{\prime}$ to launch an attack, which is much more practical for the attackers in the real world. For example, in social networks the attackers only need to sign up multiple zombie accounts and interact with real accounts. Note that, different from some injection-based attack [12, 32], NIFA will not modify the training set and true node label set $\mathcal{Y}$, as such operations are typically infeasible in the real world. The intrinsic idea of NIFA is to impact the GNN training process through massage propagation on a poisoned graph.
Within the gray-box attack settings, we also assume that the attackers have sufficient computational resources and budget to train a surrogate model and have access to the real graph as input. Similar to prior attack methods [32], attackers are also required to set thresholds $b$ and $d$ for the number of injected nodes and their degrees respectively to make NIFA deceptive and unnoticeable.